### v1

# CVPT: Cross-Attention help Visual Prompt Tuning adapt visual task

Abstract
In recent years, the rapid expansion of model sizes has led to large-scale pre-trained models demonstrating remarkable capabilities. Consequently, there has been a trend towards increasing the scale of models. However, this trend introduces significant challenges, including substantial computational costs of training and transfer to downstream tasks. To address these issues, Parameter-Efficient Fine-Tuning (PEFT) methods have been introduced. These methods optimize large-scale pre-trained models for specific tasks by fine-tuning a select group of parameters. Among these PEFT methods, adapter-based and prompt-based methods are the primary techniques. Specifically, in the field of visual fine-tuning, adapters gain prominence over prompts because of the latter’s relatively weaker performance and efficiency. Under the circumstances, we refine the widely-used Visual Prompt Tuning (VPT) method, proposing Cross Visual Prompt Tuning (CVPT). CVPT calculates cross-attention between the prompt tokens and the embedded tokens, which allows us to compute the semantic relationship between them and conduct the fine-tuning of models exactly to adapt visual tasks better. Furthermore, we introduce the weight-sharing mechanism to initialize the parameters of cross-attention, which avoids massive learnable parameters from cross-attention and enhances the representative capability of cross-attention. We conduct comprehensive testing across 25 datasets and the result indicates that CVPT significantly improves VPT’s performance and efficiency in visual tasks. For example, on the VTAB-1K benchmark, CVPT outperforms VPT over 4% in average accuracy, rivaling the advanced adapter-based methods in performance and efficiency. Our experiments confirm that prompt-based methods can achieve exceptional results in visual fine-tuning. The code is available at https://github.com/Xlgsyzp/CVPT/tree/main
1 Introduction
Increasing the scale of the models is a common method to enhance the model’s performance clip bert bart Roberta . In recent years, with the rapid development of computing devices, model sizes have significantly increased xlnet clark2020electra deberta vitg . For instance, the number of parameters in the GPT series developed by OpenAI has surged from 117 million to 1.8 trillion in just five years gpt-1 gpt-2 gpt3 . The rapidly increasing number of parameters will lead to the problem of immense computational overhead. Therefore, adapting those models to downstream tasks with the full-tuning method will incur enormous costs. To resolve this issue, the PEFT approach has been proposed lora prompt-tuning VP vl-adapter vit-adapter . PEFT adapts those large-scale pre-trained models to downstream tasks in a more efficient way by fine-tuning a subset of the models that contains much fewer parameters. Two mainstream methods within PEFT are Adapter adapter and Prompt prompt-tuning . During the training process, the Adapter inserts adapters into each transformer block and tunes those adapters, while the Prompt inserts prompt tokens into the embedded tokens to update the prompt tokens.
VPT, a prompt-based method is first introduced by Jia et al. vpt for visual fine-tuning tasks. Nevertheless, research on the adapter-based method is prominent due to its superior performance. Although some works have improved the performance of VPT dam e2vpt express , it is still challenging to match the effectiveness to that of adapter-based methods. There appears to be a consensus that prompt-based methods underperform adapter-based methods in the visual domain. But is this the case?
We conduct extensive experiments and analyses on VPT to uncover the reasons for its weaker performance compared to the Adapter. According to our experiments, we consider that the primary reason for the performance difference between VPT and adapters is that VPT’s deployment directly applies that used in NLP tasks prompt-tuning , without any adaptation to visual tasks. In NLP tasks, prompts usually contain rich semantic information that guides the fine-tuning process of the model. However, in visual tasks, prompts lack representation information. Therefore, it is necessary for VPT to use an abundant amount of prompts to fine-tune models. However, the design of VPT leads to computational inefficiency and redundancy, as well as the disruption of the self-attention between embedded tokens 3.1. As the graph follows 1, VPT shows a significant decrease in performance and an increase in costs when given a large number of prompts. Considering that, we think that VPT is unusable when given a large number of prompts.
Figure 1: Comparisons of performance and Flops between VPT and our CVPT with a pre-trained ViT-B/16 model on the VTAB-1k benchmark. We set the number of prompts to 1,10,20,50,100,150,200 respectively.
To handle the problem, we redesign VPT and introduced Cross Visual Prompt Tuning (CVPT). For the prompt tokens in CVPT, we calculate the cross-attention with the embedded tokens and add the result as residuals to the embedded tokens. This approach avoids the computational complexity of self-attention that is quadratically related to the number of prompts and allows prompts to focus on the embedded token to adapt to downstream tasks more efficiently. Additionally, by maintaining consistency in token dimensions throughout the computation process, the results of cross-attention can be directly summed with embedded tokens as residuals and do not introduce additional computational overhead for subsequent MLP. Furthermore, we share the weights of the self-attention layer with the cross-attention layer during loading checkpoints, keeping the cross-attention layer frozen alongside the self-attention layer, which eliminates the requirement for additional learned parameters for the cross-attention, and utilizes the encoded information in self-attention to help the fine-tuning of the model.
We validate the effectiveness of our method on 25 datasets, the results show that the CVPT achieves a significant improvement in performance and efficiency compared to the VPT. CVPT shows an average 4% improvement in accuracy on the 19 VTAB-1K datasets, 1% on the 5 FGVC datasets, and 3% on the ADE20K dataset. Additionally, if given fewer prompt tokens, CVPT achieves a comparable performance with other advanced PEFT methods which significantly outperforms the other prompt-based methods and needs fewer learnable parameters. If a large number of prompts is allowed, our CVPT outperforms the SOTA methods on FGVC and ADE20K datasets. Besides, although a large number of prompts are inserted, it does not introduce too much extra computational overhead compared to VPT.
Finally, we explore the impact of the deployment’s position and the effectiveness of the weight-sharing mechanism. The improvement on the model can be fully illustrated by the experimental results above, indicating that prompt-based methods can also rival SOTA adapter-based methods.
Overall, our contributions are as follows:
  • We provide a detailed analysis of the application of VPT to visual tasks, and propose that its drawback can be summarised in three points which are lack of adaptation, computational inefficiency and redundancy, destruction of self-attention.
  • We propose CVPT, which introduces cross-attention and weight-sharing mechanisms, to avoid the efficiency and performance problems caused by VPT, which allows us to use more prompts to improve performance efficiently.
  • We conducted experiments on 25 datasets with different downstream tasks. The results show that our approach significantly outperforms the original VPT and other prompt-based works in terms of performance and efficiency. It is also comparable to SOTA adapter-based methods, demonstrating the usability of the prompt-based approach for visual fine-tuning.
2 Related Work
PEFT. In the era of CNN, making bigger and deeper models was an effective way to improve performance alex resnet resnext . With the rise of transformers, this trend became even more popular. The introduction of ChatGPT further cemented the goal of the community to develop larger and more powerful models. However, limited by their scale, despite their powerful performance and generality, these large models are difficult to adapt downstream tasks by using traditional paradigms (full-tuning). Consequently, NLP researchers first proposed PEFT methods. Their works demonstrate that fine-tuning just a small number of parameters in a large-scale pre-trained model can achieve nearly the same performance as full-tuning. Encouraged by the success in NLP, researchers began to apply PEFT to large-scale vision models on different visual tasks bayesian-prompt aim . After development in the past several years, the mainstream PEFT methods can be broadly categorized into adapter-based methods and Prompt-based methods.
Adapter. Jie et al. adapter proposed inserting adapters into the network to efficiently fine-tune the model. These adapters are commonly a small network that usually contains an upsampling layer and a downsampling layer. The input is multiplied with a scaling factor $\gamma$ after passing through the upsampling and downsampling layers and then the result is added as a residual to the input. The general form of adapter can be expressed as:
$X_{out}=X_{in}+\gamma(W_{up}(W_{down}(X_{in}))),$ (1)
where $X_{in}$ denotes the input of Adapter, $\gamma$ represents the scaling factor of Adapter, and $W_{up}$ and $W_{down}$ correspond to the upsampling layer and downsampling layer, respectively. Some works did some adaption to visual tasks based on Adapter, developing several variants such as AdaptFormer adaptformer , LoRA lora and RepAdapter repAdapter etc. These adapter-based methods dominate the field of visual fine-tuning.
Prompt. Prompt was originally used in the field of NLP which is added to the input text for comprehension tasks. Lester et al. prompt-tuning proposed treating the prompt as a continuous vector and fine-tuning the model by updating its gradients. Jia et al. vpt introduced this concept to visual fine-tuning for the first time, naming it VPT. As shown in Fig.3, the embedded tokens are spliced with the prompt tokens before entering each transformer block, allowing it to participate in every layer of the network within the transformer block. Before entering the next transformer block, the prompt tokens of the previous layer are discarded, and new prompt tokens are spliced with the embedded token again (VPT-Deep). This can be formulated as shown below:
$[\vec{x}_{i},\underline{\hskip 8.5359pt},\vec{E}_{i}]$ $={\color[rgb]{0.1,0.8,0.9}L_{i}}([\vec{x}_{i-1},{\color[rgb]{0.79296875,0.046875,0.0859375}\vec{P}_{i-1}},\vec{E}_{i-1}]),$ (2)
where the red and blue indicate learnable and frozen parameters, respectively. $P$ denotes a learnable d-dimensional vector, X is the CLS token, and E is the patched image. Although there are improved variants based on VPT, such as E2VPT e2vpt , EXPRESS express and DAM-VP dam , a performance gap remains between prompt-based and adapter-based approaches.
3 Method
3.1 Analysis of previous VPT
Firstly, we analyze VPT deeply to explore why it is not better than adapter in terms of performance and efficiency, our analysis follows three points:
Lack of adaptation to visual tasks. In NLP, each token represents an actual word with rich semantic information. Therefore, the processing of concatenating prompt tokens and embedded tokens is natural and suitable for NLP tasks. However, in visual tasks, tokens represent image patches and contain sparse semantic information compared to those in NLP. Therefore, simply splicing the prompt tokens with the embedded tokens may not provide sufficient guidance information. Additionally, visual tasks often require a deeper understanding of spatial relationships and structural features of an image, which are difficult to achieve with prompt tokens.
Computational inefficiency and redundancy. When computing self-attention, the attention between each token and all other tokens needs to be calculated. Its computational complexity is $n^{2}$, where $n$ is the number of embedded tokens. If $m$ represents the number of inserted prompt tokens, the computational complexity of self-attention in VPT can be expressed as $(n+m)^{2}$. This increases the computational overhead significantly, especially when using a larger number of prompt tokens. Additionally, we found that prompt tokens are involved in the MLP computation process, which not only adds computational overhead but also does not impact the results. Our experiments show that removing the prompt token after self-attention does not affect the results.
Destruction of self-attention between embedded tokens. After softmax, the sum of the weights of all tokens is normalized to 1. Whereas, due to the addition of the prompt tokens, the sum of the weights of the embedded tokens is reduced by the prompt tokens, which corresponds to the weakening of the representation ability of the self-attention between embedded tokens. Since the prompt token is eventually removed, this is equivalent to multiplying the self-attention result between the embedded tokens by a factor which less than one. To explore how large this effect is, we set the number of prompts to 1,5,20,50,100,150,196 respectively, and visualize the tensor after the softmax function, the results are shown in Fig.2 below.
Figure 2: Self-attention weight obtained by prompt tokens and embedded tokens. We visualize the self-attention of $cls_{token}$ and exclude itself to observe the attention of $cls_{token}$ to other tokens. And the darker the color, the larger the weight. When giving 196 prompts, the attention weight obtained by prompts is over 80%, which greatly influences the self-attention received by embedded tokens.
As the number of prompts increases, the sum of the prompt’s weight values exceeds 0.8, which is over 4 times that of embedded tokens, significantly disrupting the self-attention between the embedded tokens. This explains why VPT performance decreases substantially with a larger number of prompts.
3.2 Cross Visual Prompt Tuning
Cross-Attention. Unlike self-attention vaswani2017attention , which computes the relationship between each element in the input sequence, cross-attention computes attention on two different sequences to process the semantic relationship between them crossvit . For example, in translation tasks, cross-attention is used to compute the attention weights between the source language sentence and the target language sentence. In our method, we introduce cross-attention to handle the semantic relationship between embedded tokens and prompt tokens, guiding the fine-tuning of the model. Specifically, the input of cross-attention consists of two parts: $X_{1}$ and $X_{2}$, in which $X_{1}\in\mathbb{R}^{n\times d_{1}}$ and $X_{2}\in\mathbb{R}^{m\times d_{2}}$. And $X_{1}$ serves as the query set and $X_{2}$ serves as the key-value set. We set $Q=X_{1}W^{Q}$ and $K=V=X_{2}W^{K}$, and then the cross-attention can be expressed as follows:
$CrossAttention(X_{1},X_{2})=Softmax\left(\frac{Q\cdot K}{\sqrt{d_{k}}}\right)V.$ (3)
In which $W^{Q}\in\mathbb{R}^{d_{1}\times d_{k}}$ and $W^{K}\in\mathbb{R}^{d_{2}\times d_{k}}$ are learned projection matrix, $d_{k}$ is the dimension of value-key set. In our methods, $d_{1}=d_{2}=d_{k}$. And the shape of output is $n\times d_{k}$, which is consistent with $X_{1}$.
Figure 3: Structure comparison of VPT and CVPT. In which blue represents frozen parameters and orange represents learnable parameters.
Cross Visual Prompt Tuning. We redesign the prompt to better adapt visual tasks and proposed CVPT. Our approach, as illustrated in Fig.3, follows the VPT, the main parameters of the network remain frozen, and only the final classification layer and the prompt are trainable. The key difference is that we allow the prompt token to perform cross-attention with the embedded tokens and the result of cross-attention is added with the embedded tokens as residuals. This operation helps prompts adapt visual tasks a lot, and we demonstrate how significant this improvement is in Sec.4.2. Specifically, for any input $x_{i}$ of a transformer block, the forward flow can be represented as follows:
$X_{1}=X_{i}+{\color[rgb]{0.1,0.8,0.9}SA}({\color[rgb]{0.1,0.8,0.9}LN_{1}}(X_{i})),$ (4)
$X_{2}=X_{1}+{\color[rgb]{0.1,0.8,0.9}CA}(X_{1},{\color[rgb]{0.79296875,0.046875,0.0859375}Prompt}),$ (5)
$X_{out}=X_{2}+{\color[rgb]{0.1,0.8,0.9}MLP}({\color[rgb]{0.1,0.8,0.9}LN_{2}}(X_{2})),$ (6)
where blue denotes frozen parameters and red denotes trainable parameters, SA denotes self-attention, CA denotes cross-attention, and LN denotes layer normalization.
In CVPT, we only introduce linear computational overhead associated with the number of prompt tokens. It allows CVPT to use a large number of prompt tokens to improve its performance by introducing an acceptable overhead. Furthermore, CVPT preserves the original procedure of self-attention, keeping the complete representation ability of embedded tokens. We demonstrate the improvement over VPT in terms of performance and efficiency in Sec.3.3. Finally, we set embedded tokens as query set and prompt tokens as key-value set, so that we can maintain the unity of the number of channels, allowing the result of cross-attention to be directly summed with the input as a residual term.
Weight-sharing mechanism. The utilization of cross-attention, which requires a large number of learnable parameters (usually $\geq$ 30% model’s parameter number), leads to a major challenge in computational overhead. Therefore, if the parameters of them are tunable, the computational overhead of CVPT will even rival those using full-tuning. Therefore, we introduce the weight-sharing mechanism. Due to the structure of cross-attention equals to that of self-attention, we consider that the weight of self-attention is also instructive for the fine-tuning of cross-attention. Thus, we initialize the weight of cross-attention with the parameters of self-attention when loading checkpoints. It avoids the introduction of a huge number of learnable parameters in cross-attention and keeps the efficiency of our CVPT. We explore the impact of weight-sharing in 4.3 and demonstrate that frozen cross-attention is even more effective than learnable cross-attention.
3.3 Comparison with VPT
Performance improvement. To investigate how much improvement CVPT makes and the effect of the number of prompts on performance, we use different numbers of prompt tokens and conduct experiments on VTAB-1K using VPT and CVPT, respectively. The results are shown in the following Table.1:
Table 1: Performance comparisons With VPT and CVPT on VTAB-1K benchmark of different number of prompt tokens.
Method Number 1 5 10 20 50 100 150 200
VPT 71.0 73.0 73.0 72.8 72.2 69.2 66.0 64.0
CVPT 69.5 73.5 74.0 74.1 74.3 74.5 74.6 74.8
These results show that our CVPT achieves better performance in almost every case except the number of prompts equals 1. As we analyzed in Section 3.1, VPT represents a pool absolute performance on account of the lack of adaptation to visual tasks. Besides, due to the corruption of self-attention between embedded tokens, when given a larger number of prompt tokens, VPT shows significant performance degradation or even crashes. In contrast, our CVPT avoids suffering from these problems. Additionally, its performance improves as the number of prompt tokens increases. All these results above indicate that cross-attention between prompt tokens and embedded tokens helps prompts adapting the visual tasks and instruct the model’s fine-tuning more exactly.
Efficiency improvement. To explore the improvement in efficiency of CVPT, we also recorded the amount of GPU memory occupied by VPT and CVPT during training and testing as well as the total computation of the two when conducting the above experiments, and the results are shown in Fig.4 follows:
Figure 4: The trends of training memory, testing memory, and Flops with the variation in the number of prompt tokens. Where LP represents Linear Probing which only tunes the final classifier linear. We record those data on cifar100 in VTAB-1K, the batch_size is set to 32. Pre-trained model is ViT-B/16.
It can be seen that our CVPT has made significant improvements in efficiency compared to VPT especially given a large amount of prompt tokens. Although it requires slightly more GPU memory during testing compared to full-tuning which is marginal compared to VPT. Additionally, the weight-sharing mechanism allows for targeted optimization in engineering applications, letting cross-attention and self-attention share memory, further widening the efficiency gap with VPT. Moreover, the careful design of CVPT prevents explosive growth in memory and computation as the number of prompts increases. This means we can improve the performance of CVPT by increasing the number of prompts, which is more computationally efficient than other methods.
In summary, our CVPT significantly improves the performance and efficiency of VPT by introducing cross-attention and the weight-sharing mechanism, especially given a larger number of prompts. Therefore, it allows us to introduce more prompts to the prompt-based method in an efficient manner, thus improving its performance. We will demonstrate how much this improvement is and compare it with the SOTA methods in the next section.
4 Experiment
4.1 Experimental settings
Datasets. We evaluate our CVPT on both image classification and semantic segmentation tasks to verify its effectiveness. The specific datasets involved in our work are presented in the following.
  • VTAB-1K. VTAB-1K comprises 19 datasets from different domains, classified into three main categories: the Natural group (natural images captured by standard cameras) cifar svhn caltech pets , the Specialized group (professional images captured by specialized equipment, such as medical and remote sensing images) patchcamelyon erousat , and the Structured group (synthetic images from artificial environments). Each task contains only 1,000 training samples clevr kitti matthey2017dsprites . This is a primary metric for evaluating PEFT’s performance.
  • FGVC. FGVC consists of five fine-grained visual classification benchmarks, including CUB-200-2011 CUB , NABirds birds , Oxford Flowers flower , Stanford-Dogs dogs and Stanford-Cars cars . Unlike VTAB-1K, the datasets in FGVC benchmarks are complete.
  • ADE20K. ADE20K ade20k contains more than 25,000 images and is primarily used for scene perception, parsing, segmentation, multi-object recognition, and semantic understanding. This adaptation is challenging due to the huge gap between the objectives of pretraining and downstream tasks.
Baseline. We primarily use CVPT to compare with the following methods: (1) Full-tuning, (2) Adapter and its improved variants such as LoRA, Adaptformer, RepAdapter, and SPT, and (3) VPT and its variants, including E2VPT, EXPRESS and so on.
Training. We use the ViT-Base-16 model as our main model and AdamW as our optimizer. The other settings and training strategies follow those used in VPT. To avoid extensive hyperparameter search, we only select the number of prompts from [1, 5, 10, 20] for VTAB-1K. Besides, we use single NVIDIA 3090 on VTAB-1K and FGVC benchmark, and use NVIDIA 3090 $\times$ 8 on ADE20k.
4.2 Comparison with the SOTA
VTAB-1K. We compared our method with other baseline methods on the VTAB-1K benchmark. The experimental results are shown in Table.2, where we report the top-1 accuracy of these methods. In the table, we divide the prompt-based methods into one group and the other methods into another group. The bold values in each group represent the best accuracy.
Table 2: Performance comparisons on the VTAB-1k benchmark with ViT-B/16 models pre-trained on ImageNet-21K.
Natural Specialized Structured
Method
Params. (M)
Avg. Acc.
CIFAR-100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Patch Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr/count
Clevr/distance
DMLab
KITTI/distance
dSprites/loc
dSprites/ori
SmallNORB/azi
SmallNORB/ele
Full-tuning 85.8 68.9 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1
Linear-probing mae 0 57.6 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
Bias bitfit 0.10 65.2 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1
Adapter adapter 0.15 73.9 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6
NOAH noah 0.36 75.5 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2
AdaptFormer adaptformer 0.15 74.7 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1
LoRA lora 0.29 74.5 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.4
RepAdapter repAdapter 0.23 76.1 72.4 91.6 71.0 99.2 91.4 90.7 55.1 85.3 95.9 84.6 75.9 82.3 68.0 50.4 79.9 80.4 49.2 38.6 41.0
SPT-Adapter spt 0.34 76.2 72.9 93.2 72.5 99.3 91.4 88.8 55.8 86.2 96.1 85.5 75.5 83.0 68.0 51.9 81.2 82.4 51.9 31.7 41.2
SPT-LoRA spt 0.48 76.4 73.5 93.3 72.5 99.3 91.5 87.9 55.5 85.7 96.2 85.9 75.9 84.4 67.6 52.5 82.0 81.0 51.1 30.2 41.3
VPT-shallow 0.06 67.8 77.7 86.9 62.6 97.5 87.3 74.5 51.2 78.2 92.0 75.6 72.9 50.5 58.6 40.5 67.1 68.7 36.1 20.2 34.1
VPT-Deep vpt 0.53 72.0 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8
EXPRESS express 0.98 72.9 78.0 89.6 68.8 98.7 88.9 89.1 51.9 84.8 96.2 80.9 74.2 66.5 60.4 46.5 77.6 78.0 49.5 26.1 35.3
DAM-VP dam 2.52 73.1 - - - - - - - - - - - - - - - - - - -
$E^{2}$ e2vpt 0.27 73.9 78.6 89.4 67.8 98.2 88.5 85.3 52.3 87.8 96.1 84.8 73.6 71.7 61.2 47.9 75.8 80.8 48.1 31.7 41.9
CVPT 0.10 76.2 73.0 90.0 73.8 99.2 91.2 90.0 54.4 84.0 96.5 87.2 75.7 78.4 66.7 50.4 81.0 81.5 52.6 33.4 43.3
We first compare our method with other prompt-based methods. The results of our experiments show that our method achieved the best performance among prompt-based methods in 16 out of 19 datasets, significantly outperforming VPT and other VPT-based methods. Notably, CVPT achieves the highest accuracy in all datasets within the structured group, indicating that the addition of cross-attention significantly improves the adaptation of prompts. Therefore, CVPT performs better in those out-of-distribution (OOD) datasets. Additionally, since we use fewer than 20 prompts in VTAB-1K, CVPT requires the lowest number of parameters.
When considering all PEFT methods, we find that on a small dataset like VTAB-1K, almost all mainstream PEFT methods outperformed full-tuning in terms of performance. This suggests that correctly selecting the parameters to fine-tune is crucial. For our CVPT, it shows an impressive performance, which is only 0.2% behind SPT in accuracy while using fewer parameters than SPT, and outperforms the other PEFT methods in performance. This indicates that CVPT reaches SOTA in terms of both performance and parameter count. In particular, compared to other prompt-based methods that show weaknesses, our CVPT deeply explores the potential of prompt-based methods and demonstrates that prompt-based methods can also perform well in the field of visual fine-tuning.
FGVC. Performance on VTAB-1K alone is not enough to prove the superiority of CVPT. Therefore, we introduce the experimental results of CVPT on FGVC to explore its performance on a complete dataset of a certain scale. The results are shown in Table.3 below:
Table 3: Performance comparisons on five FGVC datasets with ViT-B/16 models pre-trained on ImageNet-21K.
Method datasets  CUB-200 -2011  NABirds  Oxford Flowers  Stanford Dogs  Stanford Cars   Avg.   Acc.  Params. (M)
Full fine-tuning 87.3 82.7 98.8 89.4 84.5 88.5 86.0
Linear probing mae 85.3 75.9 97.9 86.2 51.3 79.3 0.18
Adapter  adapter 87.1 84.3 98.5 89.8 68.6 85.7 0.41
AdaptFormer  adaptformer 84.7 75.2 97.9 84.7 83.1 85.1 0.37
Bias bitfit 88.4 84.2 98.8 91.2 79.4 88.4 0.28
VPT-Shallow 86.7 78.8 98.4 90.7 68.7 84.6 0.25
VPT-Deep vpt 88.5 84.2 99.0 90.2 83.6 89.1 0.85
DAM-VP dam 87.5 82.1 99.2 92.3 - - -
EXPRESS express 88.3 - 99.0 90.0 80.5 - -
$E^{2}$ e2vpt 88.5 84.2 99.0 90.2 83.6 89.2 0.45
SPT-Adapter spt 89.1 83.3 99.2 91.1 86.2 89.8 0.41
SPT-LoRA spt 88.6 83.4 99.5 91.4 87.3 90.1 0.48
CVPT 89.7 86.1 99.3 91.4 84.9 90.3 0.79
Similar to the results on VTAB-1K, our approach substantially outperforms other prompt-based methods on FGVC benchmark. Additionally, it surpasses SPT and other adapter-based methods to achieve the best performance. This suggests that CVPT exhibits better performance on relatively large datasets like FGVC, which proves the adaptability of CVPT to the increasing scale of data in the future.
ADE20K. Finally, we apply CVPT to SETRsetr on the ADE20K dataset to explore its performance on downstream tasks of semantic segmentation. The results are shown in Table.4 below:
Table 4: Results of ADE20K datasets with ViT-L models. We report "mIoU-SS" and "mIoU-Ms" which denote single-scale and multi-scale, respectively
Methods Params(M) mIoU-SS mIoU-Ms
Full-tuning 318.3 48.31 50.07
Linear probing 13.18 35.12 37.46
Bias bitfit 13.46 43.40 45.33
VPT vpt 13.43 42.11 44.06
RepAdapter repAdapter 13.82 44.44 46.71
SPT-Adapter spt 14.60 45.20 47.20
SPT-LoRA spt 14.60 45.40 47.50
CVPT(P=10) 13.43 43.78 45.85
CVPT(P=200) 18.00 45.66 47.92
This task is quite challenging because of the huge distribution gap between pre-training datasets and downstream tasks. In this situation, our CVPT shows a 1.7% enhancement of "mIoU-SS" over the VPT with the same number of prompts. If we use 200 prompts for fine-tuning, CVPT represents a significant improvement over the other PEFT methods. This fully demonstrates the adaptation of CVPT to OOD datasets. Besides, due to our optimization of the deployment, even though the number of learnable parameters increases by 4 million, our memory usage and training time increase by less than 10% compared to linear probing and less than 5% compared to it when using 10 prompts during training.
4.3 Ablation Studies
The impact of the location of the Cross-Attention (CA). We conducted experiments with the following five positions to explore the optimal deployment of CA, and the results of the experiments are displayed in Table.5:
Position Avg. Acc.
1 73.9
2 73.9
3 74.1
4 73.3
5 73.6
Figure 5: (a) The deployments of cross-attention in ViT. Five possible positions can be inserted. Our final deployments are in dark blue. (b) Performance comparisons of different deployments of cross-attention.
We can see that inserting in prompt tokens after self-attention (SA) is the best way to perform. However, if a slight performance decrease is acceptable, we can choose position 2 to insert in parallel to improve the efficiency of the operation (this improvement is also slight).
The impact of weight-sharing between CA and SA. We set CA to be learnable (without weight-sharing) and frozen (with weight-sharing) respectively to investigate the impact of weight-sharing. The results on VTAB-1K and FGVC are shown in Table.5 below:
Table 5: Performance comparisons of learnable CA and frozen CA with weight-sharing.
Setting Learnable Para(M) VTAB-1K FGVC
Nat. Spe. Str. Avg.
learnable CA 28.4 80.1 84.8 57.8 74.2 89.4
frozen CA 0.08 80.1 84.4 57.8 74.1 90.3
We find that setting CA to tunable adds a significant number of parameters, substantially increasing computational overhead. Despite the slight performance gain it brings on VTAB-1K, it lags behind the frozen CA substantially in FGVC. Therefore, We believe that the parameters of SA are valuable for guiding the fine-tuning of CA. Especially, when dealing with a complete dataset of a certain size, such as FGVC, the weight-sharing mechanism can better utilize the pre-trained capabilities of the model, thereby improving performance.
5 Conclusion
In this paper, we explore the current mainstream prompt-based method VPT deeply and analyze the reasons why it performs poorly. Consequently, we propose a simple and effective PEFT method, CVPT, which introduces the cross-attention module to compute the cross-attention between the prompt tokens and embedded tokens thus instructing the model’s fine-tuning. What more, the weights of cross-attention are come from self-attention, avoiding introducing an enormous number of additional trainable parameters and achieving better performance. We conducted extensive experiments on 25 datasets, and the results demonstrate that CVPT achieves SOTA performance. Additionally, we conducted extensive ablation experiments on CVPT, demonstrating the impact of introducing cross-attention and weight-sharing, as well as its efficiency and performance improvements over VPT. We hope our work will inspire prompt-based PEFT methods in the future. One limitation of our work is that CVPT does not explore new strategies for the initialization of prompt tokens. In VPT, the author made a complete comparison of different initialization methods. In our work, we take the same strategy with VPT. However, we still think the optimized specific initialization method is better than the general methods VPT used. Besides, this initialization will also help us understand how prompts help the model’s fine-tuning.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See abstract, introduction, method and experiments.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See conclusion. We think a good strategy which we don’t mention in this paper can help improving the performance based on our work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We don’t think our work involves that.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See experimental settings.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: See Abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See experimental settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We follow the previous works.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See experimental settings.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We don’t think our works in relation to this.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We don’t think our work involves that.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We don’t think our work involves that.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We used publicly available datasets whose licenses allow research usage.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We don’t think our work involves that.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We don’t think our work involves that.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t think our work involves that.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.